OpenStudy (anonymous):

describe a step-by-step process to solve radical equations. List each step and provide an explanation of how to complete each step with an example. answer with 150-200 words

OpenStudy (anonymous):

Since most radicals encountered are square root (in short sqrt), I'd give example involving sqrt. Suppose the radical equation is sqrt(x-1) + 3 = 6 Steps of solving radical equations are 1) Isolate radical on one side of the equation sqrt(x-1) = 6 -3 OR, sqrt(x-1) = 3 2) Square both sides of the equation to eliminate radical (sqrt(x-1))² = 3² OR, x - 1 = 9 3) Simplify and solve as you would any equaitons x - 1 = 9 OR, x = 9+1 OR, x = 10 4) Substitute answers back into original equation to make sure that your solutions are valid Putting x = 10 in original equation, sqrt(10-1) + 3 = 6 OR, sqrt(9) = 6-3 OR, sqrt(9) = 3 Which is true. Thus, x = 10 is the solution to above radical equation.

OpenStudy (anonymous):

See lower right hand corner of this help sheet: http://www.ecalc.com/pics/math/algebra.gif